### v2

@@ I-B Motivation and Contributions
As stated in the previous subsection, in PE scenarios wherein the quality of the eavesdropping link is weaker than that of the illegitimate communication link ($R_{E}<R_{D}$), the eavesdropper will utilize some scheme (such as transmitting AN) to ensure successful eavesdropping ($R_{E}\geq R_{D}$). Moreover, in ISAC systems, the strong radar signals have been utilized as inherent jamming signals to fight the external eavesdroppers and improve the secrecy performance of the legitimate. The above works implement communication and sensing functions within the ISAC system, but the results can not be utilized in PE scenarios because of the different motivations between the PLS and PE. Thus, we have the following questions: How do we utilize the radar signals in PE scenarios? What is the effect of joint PE and TS? To answer these questions, we propose two waveform design schemes: one that maximizes the eavesdropping rate and the other that minimizes the CRB. Then, a normalized weighted optimization problem is formulated to optimize PE and TS jointly. Relative to realizing the trade-off between PLS and TS in many outstanding works, such as [7], [18], [19], [21], [22], [23], [24], etc., the results in this work demonstrate that the proposed beamforming scheme can catch two birds (PE and TS) simultaneously.
@@
    We consider the scenario in which a multi-antenna BS actively wiretaps the illegal links while sensing multiple targets. The beam transmitted by the BS for TS is also utilized as an AN to interfere with illegal links and enable successful PE. We firstly optimize the transmitting beam of the BS only to maximize the eavesdropping rate or only minimize the CRB, respectively. The semi-definite relaxation (SDR) method was employed to solve the PE and TS optimization problems after dropping the rank-one constraint.
@@
    Relative to realizing the trade-off between PLS and TS in many outstanding works, such as [7], [18], [19], [21], [22], [23], [24], the proposed scheme can effectively realize PE and TS simultaneously. The sensing signals are utilized as AN in these works to suppress the illegitimate receivers, and in theory, the more interference to the illegitimate receivers, the better the secrecy performance. There is a trade-off between communication and sensing performance in ISAC systems. However, in our work, the sensing signals need to be designed carefully since excessive interference to the illegitimate receivers will result in lower eavesdropping rate. To the best of the authors’ knowledge, no open literature addresses the joint TS and PE optimization problem. More interestingly, the results in this work demonstrated that the proposed scheme can achieve a win-win in the joint TS and PE scenario.
@@ I-C Organization
$\beta_{k}$ Normalized coefficients for target reflection complex amplitude and path attenuation of target $T_{k}$
${\bm{A}}_{r}$, ${\bm{A}}_{t}$ Steering matrix for multiple targets angles of the transmitting/receiving arrays
@@
$\bm{\beta}$ Normalized coefficients for target reflection complex amplitude and path attenuation of multiple target angles
@@
$\gamma_{s}$ Requirement between the gain of the target direction and the gain of other directions
$\alpha_{\Phi}$ Threshold of the determinant of CRB
@@ II System Model
As shown in Fig. 1, we consider a joint target sensing and proactive eavesdropping system which consists of a single-antenna illegal transmission source node $\left(S\right)$, a single-antenna destination node $\left(D\right)$, $K$ sensing targets $\left({{T_{k}},k=1,\cdots,K}\right)$, and a BS $\left(E\right)$ equipped with uniform linear array with $N_{t}$ transmitting and $N_{r}$ receiving antennas. There is an illegal link for adaptive transmission between $S$ and $D$ [32], [33], and $E$ transmits signals simultaneously to sense $T_{k}$ and wiretap the information that $S$ sends to $D$. It is assumed that the location of $S$ and $D$, the CSI of $S$-$D$, and azimuth angle of $T_{k}$ are available at $E$111 Note that these information can be obtained through the preceding detection and this assumption has been utilized in lots of literature focused on ISAC, such as [7], [23]. Explicitly, the location of $S$ can be easily estimated by its transmitted signals. The locations of $D$ and $T_{k}$ can be obtained with the proposed schemes in some outstanding works on the ISAC systems, such as [7] [12], [23], [24]. . The signals transmitted by $E$ to sense $T_{k}$ is also utilized as a controllable AN to achieve successful PE. Thus, the beam is designed to improve the accuracy of the estimation as well as the PE performance.
@@ II-A Target Sensing Model
where ${\mathbf{w}}\in\mathbb{C}^{N_{t}\times 1}$ is the transmit beamforming to be designed and ${s}$ denotes the symbols with unit power. The echo received by the radar is expressed as222 Like [14, 25], since the radar host generally has a strong signal separation capability, and a variety of methods can be used for filtering, $E$ is assumed to perfectly separate the signal (the echo from $S$) from the interference (the echo from $D$). It should be noted that this is called the worst-case scenario for PLS [37] but is the best-case scenario for PE since this assumption overestimates the eavesdropper’s decode-ability in many scenarios.
@@
where ${\bm{\theta}}=\left[{{\theta_{T_{1}}},...,{\theta_{T_{K}}}}\right]^{T}$, $\theta_{T_{k}}$ is the angle of the $k$-th target wherein the first antenna is utilized as the reference point, $k=1,\cdots,K$, ${{\mathbf{A}}_{r}}=\left[{{{\mathbf{a}}_{r}}\left({{\theta_{T_{1}}}}\right),...,{{\mathbf{a}}_{r}}\left({{\theta_{T_{K}}}}\right)}\right]$, ${{\mathbf{A}}_{t}}=\left[{{{\mathbf{a}}_{t}}\left({{\theta_{T_{1}}}}\right),...,{{\mathbf{a}}_{t}}\left({{\theta_{T_{K}}}}\right)}\right]$, ${{\mathbf{a}}_{t}}\left({{\theta_{T_{k}}}}\right)=[1,{e^{j\frac{{2\pi d}}{{\lambda}}\sin\left({\theta_{T_{k}}}\right)}},...,{e^{j\frac{{2\pi d}}{{\lambda}}\left({{N_{t}}-1}\right)\sin\left({\theta_{T_{k}}}\right)}}]^{T}$, ${{\mathbf{a}}_{r}}\left({{\theta_{T_{k}}}}\right)=[1,{e^{j\frac{{2\pi d}}{{\lambda}}\sin\left({\theta_{T_{k}}}\right)}},...,{e^{j\frac{{2\pi d}}{{\lambda}}\left({{N_{r}}-1}\right)\sin\left({\theta_{T_{k}}}\right)}}]^{T}$, $d$ is the antenna spacing, $\lambda$ denotes the wavelength, ${\bm{\beta}}={\mathrm{diag}}\left({{{\left[{{\beta_{1}},...,{\beta_{K}}}\right]}^{T}}}\right)$, $\beta_{k}={\beta_{R_{k}}}+j{\beta_{I_{k}}}$ denotes the normalized parameter which depends on the complex amplitude of the target reflection and path loss of the $S-D$ link [5], [7], [23], ${\mathbf{Z}}\sim CN\left(0,\sigma^{2}{\mathbf{I}}\right)$, and $\sigma^{2}$ is noise power. According to (2), ${\mathbf{Y}}\left({\bm{\theta}}\right)\sim CN\left({{{\mathbf{A}}_{r}}\left(\bm{\theta}\right)\bm{\beta}{\mathbf{A}}_{t}^{H}\left(\bm{\theta}\right){{\mathbf{x}}_{E}},\sigma^{2}{\mathbf{I}}}\right)$.
It must be noted that, although the distance among all the nodes, and the pre-estimated azimuth angle and the reflection amplitude are available at $E$, the aim of TS in this work is to improve the estimation performance of azimuth and reflection amplitude information through carefully design of the radar waveform. We define ${\bm{\zeta}}=\left[{{{\bm{\theta}}},{\bm{\beta}_{R}},{{{\bm{\beta}}}_{I}}}\right]$. Since $\sigma^{2}{\mathbf{I}}$ is independent of the estimator [7, 24, 38], therefore, the Fisher information matrix (FIM) for ${\bm{\zeta}}$ is expressed as
@@
where ${F_{11}}=\left({{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)$, ${F_{12}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$, ${F_{22}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$, ${{\dot{\mathbf{A}}}_{r}}=\left[{\frac{{\partial{{\bf{a}}_{r}}\left({{\theta_{{T_{1}}}}}\right)}}{{\partial{\theta_{{T_{1}}}}}},...,\frac{{\partial{{\bf{a}}_{r}}\left({{\theta_{{T_{K}}}}}\right)}}{{\partial{\theta_{{T_{K}}}}}},}\right]$, ${{\dot{\mathbf{A}}}_{t}}=\left[{\frac{{\partial{{\bf{a}}_{t}}\left({{\theta_{{T_{1}}}}}\right)}}{{\partial{\theta_{{T_{1}}}}}},...,\frac{{\partial{{\bf{a}}_{t}}\left({{\theta_{{T_{K}}}}}\right)}}{{\partial{\theta_{{T_{K}}}}}}}\right]$, and ${\mathbf{R}}={\mathbf{w}}{{\mathbf{w}}^{H}}$. The CRB matrix is obtained as
@@ II-B Proactive Eavesdropping Model
It is assumed that the link between $E$ and $D$ is line-of-sight333 In this work, the small-scale fading is not considered. For the scenarios where there is a none-line-of-sight link between the transmitter and receiver, the channel gains can be approximated by its expectation by utilizing Jensen’s inequality, like Refs. [39] and [40]. and the channel coefficient is expressed as [20]
@@
where $x_{S}$ denotes the signal sent by $S$, and ${n_{D}}\sim N\left({0,\sigma_{D}^{2}}\right)$ is the AWGN. The SINR of received signal at $D$ is expressed as
@@
where $h_{SE}=\sqrt{\frac{{{\beta_{0}}}}{{d_{SE}^{\alpha}}}}$, $d_{SE}$ represents the distance between $S$ and $E$.
@@ III-A Proactive Eavesdropping Only
It is assumed that there is an adaptive information transmission link between $S$ and $D$, $E$ can decode the received information without error, and the success of PE needs to meet ${\gamma_{D}}\leq{\gamma_{E}}$ [32, 33]. In this subsection, the transmitting beam is designed only to meet the eavesdropping demand. To maximize the eavesdropping rate on the premise of successful PE, the following optimization problem is formulated
@@
where (14a) is the constraint on total power with the maximum allowable power budget $P_{0}$ of $E$, (14b) indicates that the covariance matrix $\mathbf{R}$ is a positive semi-definite matrix, (14c) indicates that the rank of $\mathbf{R}$ is one, and (14d) is the premise of successful eavesdropping. Based on (11) and (13), (14d) is rewritten as444 It is unnecessary for $E$ to send AN to $D$ when the channel quality of $S-E$ is better than that of $S-D$, i.e. $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$ [41]. The eavesdropping rate is maximized only if the interference in the $D$’s direction is zero, in particular, $\mathcal{I}\left({{\theta_{D}}}\right)=0$. Here we mainly consider the nomal case where $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}>\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$.
@@
Moreover, it is easily found that maximizing eavesdropping rate equals minimizing $\mathcal{I}\left({{\theta_{D}}}\right)$, then, $\mathcal{P}_{1}$ is equivalent to

### v3

@@ I-B Motivation and Contributions
As stated in the previous subsection, in PE scenarios wherein the quality of the eavesdropping link is weaker than that of the illegitimate communication link ($R_{E}<R_{D}$), the eavesdropper will utilize some scheme (such as transmitting AN) to ensure successful eavesdropping ($R_{E}\geq R_{D}$). Moreover, in ISAC systems, the strong radar signals have been utilized as inherent jamming signals to fight the external eavesdroppers and improve the secrecy performance of the legitimate. The above works implement communication and sensing functions within the ISAC system, but the results can not be utilized in PE scenarios because of the different motivations between the PLS and PE. Thus, we have the following questions: How do we utilize the radar signals in PE scenarios? What is the effect of joint PE and TS? To answer these questions, we propose two waveform design schemes: one that maximizes the ER and the other that minimizes the DCRB. Then, a normalized weighted optimization problem is formulated to optimize PE and TS jointly. Relative to realizing the trade-off between PLS and TS in many outstanding works, such as [8], [20], [21], [23], [24], [25], [26], etc., the results in this work demonstrate that the proposed beamforming scheme can catch two birds (PE and TS) simultaneously.
@@
    We consider the scenario in which a multi-antenna BS actively wiretaps the illegal links while sensing multiple targets. The beam transmitted by the BS for TS is also utilized as an AN to interfere with illegal links and enable successful PE. We firstly optimize the transmitting beam of the BS only to maximize the ER or only minimize the DCRB, respectively. The semi-definite relaxation (SDR) method was employed to solve the PE and TS optimization problems after dropping the rank-one constraint.
@@
    Relative to realizing the trade-off between PLS and TS in many outstanding works, such as [8], [20], [21], [23], [24], [25], [26], the proposed scheme can effectively realize PE and TS simultaneously. The sensing signals are utilized as AN in these works to suppress the illegitimate receivers, and in theory, the more interference to the illegitimate receivers, the better the secrecy performance. There is a trade-off between communication and sensing performance in ISAC systems. However, in our work, the sensing signals need to be designed carefully since excessive interference to the illegitimate receivers will result in lower ER. To the best of the authors’ knowledge, no open literature addresses the joint TS and PE optimization problem. More interestingly, the results in this work demonstrated that the proposed scheme can achieve a win-win in the joint TS and PE scenario.
@@ I-C Organization
$\beta_{k}$
Normalized coefficients for target reflection
complex amplitude and path attenuation of target $T_{k}$
${\bm{A}}_{r}$, ${\bm{A}}_{t}$
Steering matrix for multiple targets angles
of the transmitting/receiving arrays
@@
$\bm{\beta}$
Normalized coefficients for target reflection complex
amplitude and path attenuation of multiple target angles
@@
$\gamma_{s}$
Requirement between the gain of
the target direction and the gain of other directions
$\alpha_{\Phi}$ Threshold of the DCRB
@@ II System Model
As shown in Fig. 1, we consider a joint TS and PE system which consists of a single-antenna illegal transmission source node $\left(S\right)$, a single-antenna destination node $\left(D\right)$, $K$ sensing targets $\left({{T_{k}},k=1,\cdots,K}\right)$, and a BS $\left(E\right)$ equipped with uniform linear array with $N_{t}$ transmitting and $N_{r}$ receiving antennas. There is an illegal link for adaptive transmission between $S$ and $D$ [35], [36], and $E$ transmits signals simultaneously to sense $T_{k}$ and wiretap the information that $S$ sends to $D$. It is assumed that the location of $S$ and $D$, the CSI of $S$-$D$ link, and azimuth angle of $T_{k}$ are available at EE111 Note that these information can be obtained through the preceding detection and this assumption has been utilized in lots of literature focused on ISAC, such as [8], [25]. Explicitly, the location of $S$ can be easily estimated by its transmitted signals. The locations of $D$ and $T_{k}$ can be obtained with the proposed schemes in some outstanding works on the ISAC systems, such as [8] [14], [25], [26]. The parameter estimation consists of rough estimation, waveform design, and fine estimation, which form a closed-loop optimization logic. This work focuses on waveform design, which serves as a “bridge” connecting rough and fine estimation. Based on the rough estimation results, a transmission waveform that is insensitive to errors is designed to enhance the recognizability of the target parameters in the received signal. . The signals transmitted by $E$ to sense $T_{k}$ is also utilized as a controllable AN to achieve successful PE. Thus, the beam is designed to improve the accuracy of the estimation as well as the PE performance.
@@ II-A Target Sensing Model
where ${\mathbf{w}}\in\mathbb{C}^{N_{t}\times 1}$ is the transmit beamforming to be designed and s{s} denotes the symbols with unit power. The echo received by the radar is expressed as222 Like [16, 27], since the radar host generally has a strong signal separation capability, and a variety of methods can be used for filtering, $E$ is assumed to perfectly separate the signal (the echo from $S$) from the interference (the echo from $D$). It should be noted that this is called the worst-case scenario for PLS [40] but is the best-case scenario for PE since this assumption overestimates the eavesdropper’s decode-ability in many scenarios.
@@
where ${\bm{\theta}}=\left[{{\theta_{T_{1}}},...,{\theta_{T_{K}}}}\right]^{T}$, $\theta_{T_{k}}$ is the angle of the $k$-th target wherein the first antenna is utilized as the reference point, $k=1,\cdots,K$, ${{\mathbf{A}}_{r}}=\left[{{{\mathbf{a}}_{r}}\left({{\theta_{T_{1}}}}\right),...,{{\mathbf{a}}_{r}}\left({{\theta_{T_{K}}}}\right)}\right]$, ${{\mathbf{A}}_{t}}=\left[{{{\mathbf{a}}_{t}}\left({{\theta_{T_{1}}}}\right),...,{{\mathbf{a}}_{t}}\left({{\theta_{T_{K}}}}\right)}\right]$, ${{\mathbf{a}}_{t}}\left({{\theta_{T_{k}}}}\right)=[1,{e^{j\frac{{2\pi d}}{{\lambda}}\sin\left({\theta_{T_{k}}}\right)}},...,{e^{j\frac{{2\pi d}}{{\lambda}}\left({{N_{t}}-1}\right)\sin\left({\theta_{T_{k}}}\right)}}]^{T}$, ${{\mathbf{a}}_{r}}\left({{\theta_{T_{k}}}}\right)=[1,{e^{j\frac{{2\pi d}}{{\lambda}}\sin\left({\theta_{T_{k}}}\right)}},...,{e^{j\frac{{2\pi d}}{{\lambda}}\left({{N_{r}}-1}\right)\sin\left({\theta_{T_{k}}}\right)}}]^{T}$, $d$ is the antenna spacing, $\lambda$ denotes the wavelength, ${\bm{\beta}}={\mathrm{diag}}\left({{{\left[{{\beta_{1}},...,{\beta_{K}}}\right]}^{T}}}\right)$, $\beta_{k}={\beta_{R_{k}}}+j{\beta_{I_{k}}}$ denotes the normalized parameter which depends on the complex amplitude of the target reflection and path loss of the SDS-D link [6], [8], [25], ${\mathbf{Z}}\sim CN\left(0,\sigma^{2}{\mathbf{I}}\right)$, and $\sigma^{2}$ is noise power. According to (2), ${\mathbf{Y}}\left({\bm{\theta}}\right)\sim CN\left({{{\mathbf{A}}_{r}}\left(\bm{\theta}\right)\bm{\beta}{\mathbf{A}}_{t}^{H}\left(\bm{\theta}\right){{\mathbf{x}}_{E}},\sigma^{2}{\mathbf{I}}}\right)$.
It must be noted that, although the distance among all the nodes, and the pre-estimated azimuth angle and the reflection amplitude are available at $E$, the aim of TS in this work is to improve the estimation performance of azimuth and reflection amplitude information through careful design of the radar waveform. We define ${\bm{\zeta}}=\left[{{{\bm{\theta}}},{\bm{\beta}_{R}},{{{\bm{\beta}}}_{I}}}\right]$. Since $\sigma^{2}{\mathbf{I}}$ is independent of the estimator [8, 26, 41], therefore, the Fisher information matrix (FIM) for $\bm{\zeta}$} is expressed as
@@
where ${F_{11}}=\left({{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)$, ${F_{12}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$, ${F_{22}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$, ${{\dot{\mathbf{A}}}_{r}}=\left[{\frac{{\partial{{\bf{a}}_{r}}\left({{\theta_{{T_{1}}}}}\right)}}{{\partial{\theta_{{T_{1}}}}}},...,\frac{{\partial{{\bf{a}}_{r}}\left({{\theta_{{T_{K}}}}}\right)}}{{\partial{\theta_{{T_{K}}}}}},}\right]$, ${{\dot{\mathbf{A}}}_{t}}=\left[{\frac{{\partial{{\bf{a}}_{t}}\left({{\theta_{{T_{1}}}}}\right)}}{{\partial{\theta_{{T_{1}}}}}},...,\frac{{\partial{{\bf{a}}_{t}}\left({{\theta_{{T_{K}}}}}\right)}}{{\partial{\theta_{{T_{K}}}}}}}\right]$, and ${\mathbf{R}}={\mathbf{w}}{{\mathbf{w}}^{H}}$.
Proof.
Proof: See the Appendix A. ∎
The CRB matrix is obtained as
@@ II-B Proactive Eavesdropping Model
It is assumed that the link between $E$ and $D$ is line-of-sight333 This work assumes the link between $E$ and $D$ is line-of-sight and the CSI is perfectly known. It is easy to extend the results to scenarios with small-scale fading and/or with imperfect CSI. For the scenarios where there is a non-line-of-sight link between the transmitter and receiver, the channel gains can be approximated by their expectation by utilizing Jensen’s inequality, like Refs. [29], [42] and [43]. For the bounded CSI error model, the distance between $E$ and $T_{k}$/$D$ can be approximated by its upper bound, which can be obtained based on the triangle inequality [43]. For the Gaussian CSI error model, the distances can be approximated by its expectation. and the channel coefficient is expressed as [22]
@@
where $x_{S}$ denotes the signal sent by $S$, and ${n_{D}}\sim N\left({0,\sigma_{D}^{2}}\right)$ is the Additive White Gaussian Noise. The SINR of received signal at $D$ is expressed as
@@
where $h_{SE}=\sqrt{\frac{{{\beta_{0}}}}{{d_{SE}^{\alpha}}}}$ and $d_{SE}$ represents the distance between $S$ and $E$.
@@ III-A Proactive Eavesdropping Only
It is assumed that there is an adaptive information transmission link between $S$ and $D$, $E$ can decode the received information without error, and the success of PE needs to meet ${\gamma_{D}}\leq{\gamma_{E}}$ [35, 36]. In this subsection, the transmitting beam is designed only to meet the eavesdropping demand. To maximize the ER on the premise of successful PE, the following optimization problem is formulated
@@
where (14a) is the constraint on total power with the maximum allowable power budget $P_{0}$ of $E$, (14b) indicates that the covariance matrix $\mathbf{R}$ is a positive semi-definite matrix, (14c) indicates that the rank of $\mathbf{R}$ is one, and (14d) is the premise of successful eavesdropping. Based on (11) and (13), (14d) is rewritten as444 It is unnecessary for $E$ to send AN to $D$ when the channel quality of SES-E is better than that of SDS-D, i.e. $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$ [44]. The ER is maximized only if the interference in the $D$’s direction is zero, in particular, $\mathcal{I}\left({{\theta_{D}}}\right)=0$. Here we mainly consider the nomal case where $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}>\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$.
@@
Moreover, it is easily found that maximizing ER equals minimizing $\mathcal{I}\left({{\theta_{D}}}\right)$, then, $\mathcal{P}_{1}$ is equivalent to